OpenStudy (amtran_bus):

Prove .99999999999999999999999999999999999999999999999999999 equals 1.

OpenStudy (anonymous):

that is false.....the number mentioned above ROUNDS UP to 1....meaning it is closer to one than it is to 0 :)

OpenStudy (amtran_bus):

Good job! Now try .9 repeating!

OpenStudy (anonymous):

its the exact same explanation thou stubborn being

OpenStudy (asnaseer):

If you mean 0.999... (i.e. 9's repeating for ever) then this can be proved

OpenStudy (amtran_bus):

Go fir it!

OpenStudy (asnaseer):

is this a challenge or are you looking for help in trying to prove it?

OpenStudy (amtran_bus):

I want to prove it.

OpenStudy (asnaseer):

ok, start by setting:\[x=0.999..\]Then see what happens if you multiply both sides by 10

OpenStudy (amtran_bus):

10x=0.999(10)

OpenStudy (amtran_bus):

or 9.999999999---

OpenStudy (amtran_bus):

10=9.99999999 on and on

OpenStudy (asnaseer):

correct - and now notice that 9.99... can be written as 9 + 0.999...

OpenStudy (amtran_bus):

right

OpenStudy (asnaseer):

and what is 0.999.... (look at how we started the proof)

OpenStudy (amtran_bus):

you mean =x?

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so you end up with: x = 0.999... 10x = 9.999... = 9 + 0.999... = 9 + x

OpenStudy (amtran_bus):

ohhhhh. Thanks!!!

OpenStudy (asnaseer):

yw :)

OpenStudy (phi):

how about x=0.9999... 10x= 9.9999... 10x-x = 9 9x=9 x=1

OpenStudy (asnaseer):

thats basically exactly what we have done :)

OpenStudy (phi):

yes, just wanted it all in one post for other readers